# Infrared and visible Image Fusion with Language-driven Loss in CLIP Embedding Space

### v1

@@ 1 Introduction
Multi-modal image fusion techniques, particularly the ones for infrared and visible image fusion (IVIF), have received considerable attention in recent yearsZhang and Demiris (2023). Infrared images reflect thermal radiation information, unaffected by lighting conditions, but often lack sufficient details. Conversely, visible images are often rich in details but sensitive to illumination, smoke and adverse weather. IVIF task aims to generate a fused image that integrate the complementary information from the two modalities. Currently, the deep-learning based methods dominate the field of IVIF. However, since there is no physical ground-truth fused image, the fusion output is greatly dependent on the loss functions defined mathematically for modelling the fused image.
@@
Researchers have proposed to use various loss functions to supervise fusion network training, including contrastive lossZhu et al. (2022), adversarial lossRao et al. (2023); Ma et al. (2020), high-level task supervision lossXie et al. (2023); Tang et al. (2023), SSIM lossLi and Wu (2018); Li et al. (2023), etc. To improve the fusion results, current deep learning-based methods tend to combine various loss functions. Consequently, the overall loss function becomes increasingly complex. However, without knowing the real fused image, it is difficult for us to really well model the fused output mathematically with analytical loss functions. As a result, the performance of these fusion methods is limited.
@@
More specifically, to achieve the above goal, we try to leverage the currently popular pre-trained vision-language model CLIPRadford et al. (2021). As shown in Figure 1, we found that CLIP is able to perceive the infrared images, indicating that it can well encode the infrared image information (as that for the typical visible image) into the multi-modal embedding space of CLIP, in which the vision-language alignment has been achieved. To solve the issue of IVIF, we next establish an embedded language-driven fusion model using the text encoder of CLIP, for which we developed a language-expressed fusion objective, and use the prompts “an infrared image” and “a visible gray image” to describe the input infrared and visible images, respectively. Our goal is then become to make the fusion of infrared and visible images aligned with the language-driven fusion model in the CLIP embedding space through supervised training.
@@ 2 Related Work
Infrared-visible Image Fusion Traditional IVIF methods utilize manually designed fusion rules and features to fuse infrared and visible information. These include multi-scale decomposition methods Vishwakarma and Bhuyan (2018); Zhou et al. (2016, 2023), sparse representation Wang et al. (2014), and saliency-based methods Han et al. (2013a); Ellmauthaler et al. (2012). In recent years, deep learning-based methods Ma et al. (2022); Cheng et al. (2023); Liu et al. (2021); Zhang et al. (2020) have become the mainstream approaches to address the inherent challenges of the IVIF task. Due to the lack of ground truth, the fusion result is highly dependent on the used loss functions. Approaches like U2Fusion Xu et al. (2020), DenseFuse Li and Wu (2018) and NestFuse Li et al. (2020) mainly utilize the pixel intensity loss and structure similarity Ma et al. (2015) loss to ensure that fusion results preserve the source image information. To obtain the fusion results with richer details, perceptual loss Long et al. (2021) are introduced for training. In addition, GAN-based methods Rao et al. (2023); Ma et al. (2020) use adversarial loss to preserve the salient thermal targets and texture details. More recent studies utilize high-level tasks, such as object detection Liu et al. (2022); Zhao et al. (2023), image classification Tang et al. (2022), and semantic segmentation Xie et al. (2023); Tang et al. (2023), to construct the semantic losses to improve the fusion performance. Although can help to obtain better object detection or segmentation performance, the fusion performance of these methods are constrained by a limited number of object or scene labels, and can not obtain a very comprehensive fused image.
Text-driven Image Generation with CLIP OpenAI proposed CLIP Radford et al. (2021), a high-performance text-image embedding model trained on 400 million text-image pairs. This model bridge the gap between image and language, establishing new benchmarks in related fields, including image generation. Techniques for generating and manipulating images driven by text increasingly utilize the CLIP model. Representative methods Patashnik et al. (2021); Gal et al. (2022) combines CLIP with StyleGAN Karras et al. (2019, 2020) to synthesize images by optimizing the latent code of StyleGAN, realizing image manipulation through text prompts. DiffusionCLIP Kim and Ye (2021) incorporates CLIP with a diffusion model Song and Ermon (2019) to enable text-baesd image manipulation, providing enhanced performance and more stable training compared with GAN-based approaches. So far, there is no literature to use CLIP to solve the challenge of multi-modal image fusion. For this purpose, a proper language-based fusion model should be established based on CLIP.
@@ 3.2 Language-expressed Fusion Objective
The CLIP model that can perceive both infrared and visible image information allows us to express the fusion objective using language, and then encode it into the CLIP embedding space. As illustrated in Figure 2, the language expression stems from our mind imaging what the fusion result should be. Using the language, it is possible for us to give a comprehensive description of the fusion effect with a short sentence, e.g., in this paper we describe it as “a vivid image with detailed background and obvious objects”. This expression can lead to produce the fused image with pretty rich and comprehensive information, covering both the background and various objects in the image. In contrast, it is difficult to use analytical loss functions to formulate this objective, because we can not define the “background” and “objects” with mathematical symbols, while the pre-trained language-vision model might have this cognition for an image. Some work Liu et al. (2022); Tang et al. (2023) leveraged the object detection or semantic segmentation results to constructed the supervision losses for IVIF. However, these approaches are constrained by a limited number of object or scene labels, whereas CLIP have the perception for various objects and backgrounds, not to mention that it can react to the words “obvious” and “detailed” in the objective description.
@@ 3.3 Language-driven Fusion Model and Loss
Embedded Language-driven Fusion Model. After having the language-expressed fusion objective encoded in the CLIP embedding space, the next issue is to establish a language-driven fusion model, which can guide to generate the described target image from the source images. As shown in Figure 3, let $V_{f}$ denote the embedded vector from the text to express the fusion objective, and $V_{vs}$, $V_{ir}$ denote the embedded vectors from the texts to describe the visible and infrared images, respectively. Inspired by the image generator with CLIP Gal et al. (2022), we construct the fusion model by establishing the transition relationship from the source vectors $V_{vs}$, $V_{ir}$ to the objective vector $V_{f}$. This results in the transition vectors $\Delta V_{vs}=V_{f}-V_{vs}$ and $\Delta V_{ir}=V_{f}-V_{ir}$ as illustrated by the dash arrows in Figure 3. The combination of these two vectors jointly defines the direction to generate the fused result from the input modalities in the embedding space.
@@ 4 Experiments
Datasets. We conduct experiments on three public datasets: M3FD Liu et al. (2022), RoadScene Xu et al. (2020), and TNO Toet (2017). We use M3FD for network training, as it covers various scenes, illumination, and weather conditions. The TNO dataset includes multi-spectral, military-relevant images, while RoadScene focuses on outdoor scenes. The TNO and RoadScene datasets are used to evaluate the performance of our proposed method.
Compared Methods. We compare our method with nine state-of-the-art fusion methods, including FusionGANMa et al. (2019), MFEIFLiu et al. (2021), PIAFusionTang et al. (2022), PMGIZhang et al. (2020), RFNLi et al. (2021), SwinFusionMa et al. (2022), U2FusionXu et al. (2020), UMFDi et al. (2022) and GANMcCMa et al. (2021).
Metrics. Five metrics are used to evaluate our method: information entropy (EN)Roberts et al. (2008), average gradient (AG)Cui et al. (2015), standard deviation (SD)Ma et al. (2019), spatial frequency (SF)Roberts et al. (2008) and the visual information fidelity for fusion (VIFF)Han et al. (2013b). EN measures the amount of information contained in the fusion result. AG measures image gradients, with higher values signifying richer detail. SD is used to quantify the contrast in images. SF evaluates the level of detail in an image. VIFF evaluates the quality of image fusion, considering how human vision perceives these images.

### v2

@@ 1 Introduction
Multi-modal image fusion techniques, particularly the ones for infrared and visible image fusion (IVIF), have received considerable attention in recent years[?]. Infrared images reflect thermal radiation information, unaffected by lighting conditions, but often lack sufficient details. Conversely, visible images are often rich in details but sensitive to illumination, smoke and adverse weather. IVIF task aims to generate a fused image that integrate the complementary information from the two modalities. Currently, the deep-learning based methods dominate the field of IVIF. However, since there is no physical ground-truth fused image, the fusion output is greatly dependent on the loss functions defined mathematically for modelling the fused image.
@@
Researchers have proposed to use various loss functions to supervise fusion network training, including contrastive loss[?], adversarial loss[??], high-level task supervision loss[??], SSIM loss[??], etc. To improve the fusion results, current deep learning-based methods tend to combine various loss functions. Consequently, the overall loss function becomes increasingly complex. However, without knowing the real fused image, it is difficult for us to really well model the fused output mathematically with analytical loss functions. As a result, the performance of these fusion methods is limited.
@@
More specifically, to achieve the above goal, we try to leverage the currently popular pre-trained vision-language model CLIP[?]. As shown in Figure 1, we found that CLIP is able to perceive the infrared images, indicating that it can well encode the infrared image information (as that for the typical visible image) into the multi-modal embedding space of CLIP, in which the vision-language alignment has been achieved. To solve the issue of IVIF, we next establish an embedded language-driven fusion model using the text encoder of CLIP, for which we developed a language-expressed fusion objective, and use the prompts “an infrared image” and “a visible gray image” to describe the input infrared and visible images, respectively. Our goal is then become to make the fusion of infrared and visible images aligned with the language-driven fusion model in the CLIP embedding space through supervised training.
@@ 2 Related Work
Infrared-visible Image Fusion Traditional IVIF methods utilize manually designed fusion rules and features to fuse infrared and visible information. These include multi-scale decomposition methods [???], sparse representation [?], and saliency-based methods [??]. In recent years, deep learning-based methods [????] have become the mainstream approaches to address the inherent challenges of the IVIF task. Due to the lack of ground truth, the fusion result is highly dependent on the used loss functions. Approaches like U2Fusion [?], DenseFuse [?] and NestFuse [?] mainly utilize the pixel intensity loss and structure similarity [?] loss to ensure that fusion results preserve the source image information. To obtain the fusion results with richer details, perceptual loss [?] are introduced for training. In addition, GAN-based methods [??] use adversarial loss to preserve the salient thermal targets and texture details. More recent studies utilize high-level tasks, such as object detection [??], image classification [?], and semantic segmentation [??], to construct the semantic losses to improve the fusion performance. Although can help to obtain better object detection or segmentation performance, the fusion performance of these methods are constrained by a limited number of object or scene labels, and can not obtain a very comprehensive fused image.
Text-driven Image Generation with CLIP OpenAI proposed CLIP [?], a high-performance text-image embedding model trained on 400 million text-image pairs. This model bridge the gap between image and language, establishing new benchmarks in related fields, including image generation. Techniques for generating and manipulating images driven by text increasingly utilize the CLIP model. Representative methods [??] combines CLIP with StyleGAN [??] to synthesize images by optimizing the latent code of StyleGAN, realizing image manipulation through text prompts. DiffusionCLIP [?] incorporates CLIP with a diffusion model [?] to enable text-baesd image manipulation, providing enhanced performance and more stable training compared with GAN-based approaches. So far, there is no literature to use CLIP to solve the challenge of multi-modal image fusion. For this purpose, a proper language-based fusion model should be established based on CLIP.
@@ 3.2 Language-expressed Fusion Objective
The CLIP model that can perceive both infrared and visible image information allows us to express the fusion objective using language, and then encode it into the CLIP embedding space. As illustrated in Figure 2, the language expression stems from our mind imaging what the fusion result should be. Using the language, it is possible for us to give a comprehensive description of the fusion effect with a short sentence, e.g., in this paper we describe it as “a vivid image with detailed background and obvious objects”. This expression can lead to produce the fused image with pretty rich and comprehensive information, covering both the background and various objects in the image. In contrast, it is difficult to use analytical loss functions to formulate this objective, because we can not define the “background” and “objects” with mathematical symbols, while the pre-trained language-vision model might have this cognition for an image. Some work [??] leveraged the object detection or semantic segmentation results to constructed the supervision losses for IVIF. However, these approaches are constrained by a limited number of object or scene labels, whereas CLIP have the perception for various objects and backgrounds, not to mention that it can react to the words “obvious” and “detailed” in the objective description.
@@ 3.3 Language-driven Fusion Model and Loss
Embedded Language-driven Fusion Model. After having the language-expressed fusion objective encoded in the CLIP embedding space, the next issue is to establish a language-driven fusion model, which can guide to generate the described target image from the source images. As shown in Figure 3, let $V_{f}$ denote the embedded vector from the text to express the fusion objective, and $V_{vs}$, $V_{ir}$ denote the embedded vectors from the texts to describe the visible and infrared images, respectively. Inspired by the image generator with CLIP [?], we construct the fusion model by establishing the transition relationship from the source vectors $V_{vs}$, $V_{ir}$ to the objective vector $V_{f}$. This results in the transition vectors $\Delta V_{vs}=V_{f}-V_{vs}$ and $\Delta V_{ir}=V_{f}-V_{ir}$ as illustrated by the dash arrows in Figure 3. The combination of these two vectors jointly defines the direction to generate the fused result from the input modalities in the embedding space.
@@ 4 Experiments
Datasets. We conduct experiments on three public datasets: M3FD [?], RoadScene [?], and TNO [?]. We use M3FD for network training, as it covers various scenes, illumination, and weather conditions. The TNO dataset includes multi-spectral, military-relevant images, while RoadScene focuses on outdoor scenes. The TNO and RoadScene datasets are used to evaluate the performance of our proposed method.
Compared Methods. We compare our method with nine state-of-the-art fusion methods, including FusionGAN[?], MFEIF[?], PIAFusion[?], PMGI[?], RFN[?], SwinFusion[?], U2Fusion[?], UMF[?] and GANMcC[?].
Metrics. Five metrics are used to evaluate our method: information entropy (EN)[?], average gradient (AG)[?], standard deviation (SD)[?], spatial frequency (SF)[?] and the visual information fidelity for fusion (VIFF)[?]. EN measures the amount of information contained in the fusion result. AG measures image gradients, with higher values signifying richer detail. SD is used to quantify the contrast in images. SF evaluates the level of detail in an image. VIFF evaluates the quality of image fusion, considering how human vision perceives these images.